OpenStudy (anonymous):

Mercury boils at 357°C and its average enthalpy of vaporization is 60.9 kJ/mol. Estimate the equilibrium vapour pressure of mercury at 25°C.

OpenStudy (anonymous):

Hey @ Jadedry, I don't know if you're working on this or not but I appreciate the help. I know this is a Clausius-Clapeyron Equation type question and I've set it up as such but its the inital pressure I can't figure out what to put for it, or maybe I don't put anything since its the "equilibrium vapour pressure"

OpenStudy (jadedry):

I'm typing it out! just bear with me. c:

OpenStudy (jadedry):

If you're just worried about what to put for initial pressure you should try 760 torr.

OpenStudy (anonymous):

alright, i know its not easy typing these questions

OpenStudy (jadedry):

Which is the pressure at 1 atmosphere, it seems like the best option if they don't tell you what pressure they're working at. (Yeah, it does take a while, sorry! ;n;)

OpenStudy (jadedry):

You have to use the Clausius-Clapeyron Equation. which is: \[\ln(P _{1} / P _{2}) = -(\Delta H/R) (\frac{ 1 }{ T_{1} } - \frac{ 1 }{ T _{2}})\] Where P1 is the equilibrium vapour pressure of mercury at 25°C, and P2 is 760 torr (this is the value of torr at 1 atmosphere; since the pressure is not stated, you use 1 atmosphere.) Delta H is average enthalpy of vaporization (in joules!) and R is the gas constant, which equals: 8.31447 joules/kelvin/mole T1 and T2 = your given temperatures. You should (I think) end up with this: \[\ln(P _{1}/760) = - (\frac{ 60900 }{ 8.31447 })(\frac{ 1 }{ 298.15 }-\frac{ 1 }{ 630.15 })\]

OpenStudy (jadedry):

And then solve for P1. (I converted the temps to kelvin.)

OpenStudy (anonymous):

Ya this was my exact set-up except i didnt use torr

OpenStudy (jadedry):

if you want atmosphere you could just use 1 then.

OpenStudy (jadedry):

in place of 760.

OpenStudy (anonymous):

This was more of a second opinion question aha. I have another if you have the time i'll close this one so you can gain another medal

OpenStudy (jadedry):

Thanks! I'll have a look at the next one you post and see if I can help. ;u;